OpenStudy (anonymous):

Let f^(n) (a) denote the n-th derivative at a. If f(x)=sin(x)cos(2x)+cos(x)sin(2x) find f^(52) (π/18).

OpenStudy (anonymous):

f(x)=sin 3x f'(x)=3cos 3x=\[3\sin \left( \frac{ \pi }{ 2 }+3x \right)\] \[f"(x)=3^{2}\cos \left( \frac{ \pi }{2 }+3x \right)=3^{2}\sin \left( \frac{ \pi }{ 2 } +\frac{ \pi }{2 }+3x\right)=3^{2}\sin \left( 2*\frac{ \pi }{2 }+3x \right)\] \[f \prime \prime \prime \left( x \right)=3^{3}\cos \left( 2*\frac{ \pi }{2 } +3x\right)=3^{3}\sin \left( \frac{ \pi }{2 }+2*\frac{ \pi }{2 }+3x \right)=3^{3}\sin \left( \frac{ 3*\pi }{2}+3x \right)\] ............................................................................................................................. .............................................................................................................................. ............................................................................................................................. \[f ^{n}\left( x \right)=3^{n}\sin \left( n*\frac{ \pi }{2 }+3x \right)\] put n=52, \[x=\frac{ \pi }{18}and get the result.\]

OpenStudy (anonymous):

thank you!